The perimeter of a rectangle is 80 ft. Find the dimensions of the length is 5 feet longer than 4 times the width. What is the area of the rectangle?
Answer by stanbon(75887)

The perimeter of a rectangle is 80 ft. Find the dimensions if the length is 5 feet longer than 4 times the width. What is the area of the rectangle?
-------------------
Let the width be "x" ft.
Then length = 4x+5 ft
-------------------
EQUATION:
Perimeter = 2(length + width)
80 = 2(4x+5+x)
40 = 5x+5
x = 7 ft
4x+5 = 4*7+5 = 33 ft
-----------------
Area = 7*33 = 231 sq. ft.
=============
Cheers,
Stan H.